(c) Steam flows a pipe of external diameter 30 mm, where the pipe outer surface temperature is 130oC. Circular annular fins (k = 180 W·m-1·K-1) of outer diameter 60 mm and thickness 1 mm are fitted tightly to the pipe. Heat is transferred by convection (h=10 W·m-2·K-1) to the surrounding air at 20oC.

  1. (i) Determine the efficiency of a single fin.

  2. (ii) Determine the heat transfer rate for a single fin.

(8 marks) (6 marks)

(iii) Determine the effectiveness of a 1 m length of finned pipe, assuming a 4 mm gap between

successive fins.
Answer #1

(C)

(i)

The efficiency of a single fin:

\eta = 97\%

(ii)

The heat transfer rate for a single fin:

\dot{Q}_{fin}=4.725W

(iii)

The effectiveness of a 1 m length of finned pipe:

\epsilon =9.92

The detailed solution is given below:

Giveni> 30mm Tp = 130°c Too = 20°C lo w m?.K h = =Amm 18ow/m.k ivey =7 d = 30mm 1o0 (Thi knes) D = 60mm d = 30mm = 0.015m = 0

low/m?.k 180w/m.k Xx 0.00m (0:015 +x 0.001X 0.1155 and Det = 2.03,We get # Noww for g - 0.1155 → Fin efkiacy, , - 0.97 = 170

4.725 W [Le lm] 1 gah,g = m (00 No fin TA(0.03)x| 0.0942m? Ang fon hx Ano, fin X (To-To) 0.0942m² }x ( Bo+273)-(20+273) īdL 1

4mm Imm< Amm 7. Imm lomm From figure we (on see that in lonm Number of in = Vum ber of Pm ter fin Surtace = 2 In Im ling th,